OpenStudy (anonymous):

Points P=(1,0,0), Q=(0,2,0), R=(0,0,3). Is the vector PQ orthagonal to the vector QR why or why not?

OpenStudy (amistre64):

create the 2 vectors and dot them

OpenStudy (anonymous):

Find an equation of the plane containing P, Q, and R.

OpenStudy (amistre64):

cross the 2 vectors from the first question you asked to determine a "normal" vector then create the equation using the normal vector and any of the given points

OpenStudy (anonymous):

\[\rightarrow PR X rightarrowPQ \]

OpenStudy (anonymous):

I'm trying to show you here what I have on my paper..

OpenStudy (amistre64):

that looks like its going in the right direction; i tend to use a different setup tho first lets check that your vectors are good P=(1,0,0), Q=(0,2,0) -1 -1 ----------------- 0 0 0 -1 2 0 Q=(0,2,0), R=(0,0,3) -2 -2 ----------------- 0 0 0 0 -2 3 i see you used PR instead of PQ ... which is fine, any 2 vectors from the set will work

OpenStudy (amistre64):

i tend to do a column setup instead of a row setup; and i think that QR will be easier for me to play with x -1 0 ; x = 6 y 2 -2 ; -y = -3 z 0 3 ; z = 2

OpenStudy (amistre64):

of course your vectors produce the same results x -1 -1 ; x = 6 y 2 0 ; -y = -3 z 0 3 ; z = 2

OpenStudy (amistre64):

normal (6,3,2,), attach this to a given point (a,b,c) to produce 6(x-a) + 3(y-b) + 2(z-c) = 0

OpenStudy (anonymous):

I have attached a document showing my work and final answer please let me know if it is correct or of any errors I made. Thank you.

OpenStudy (amistre64):

your cross is alittle off, and once you find the crossing vector, you have to apply to develop a dot product to all vectors of a given point.

OpenStudy (amistre64):

spose we have a point: (xo,yo,zo) all vectors from this point to any given (x,y,z) point is (x-xo, y-yo, z-zo) to define all the vectors of this set that are perpendicular to a given vector (a,b,c), we do a dot product that results in a(x-xo) + b(y-yo) + c(z-zo) = 0

OpenStudy (amistre64):

the crossing of 2 vectors will produce the (a,b,c) vector we need to define the plane, and any given point can be used as the (xo,yo,zo) to anchor it to

OpenStudy (amistre64):

a cross vector is only as good as the method employed to produce it. your setup has a few flaws in it

OpenStudy (amistre64):

\[\begin{vmatrix}2&0\\0&3\end{vmatrix}i-\begin{vmatrix}-1&0\\-1&3\end{vmatrix}j+\begin{vmatrix}-1&2\\-1&0\end{vmatrix}k\] \[6i-(-3)j+2k\] \[6i+3j+2k\to~<6,3,2>\]

OpenStudy (anonymous):

so 6i + 3j+ 2k would be my final answer?

OpenStudy (amistre64):

of course not, that is only the normal vector of the plane equation.

OpenStudy (amistre64):

you need to determine which point you want to use to fill in the actual plane equation with: P Q or R ?? this will be the values of (xo,yo,zo) in the setup 6(x-xo) + 3(y-yo) + 2(z-zo) = 0

OpenStudy (anonymous):

=( I'm confused I thought I was on the right track...

OpenStudy (amistre64):

on the right track, yes at the end of the track, no

OpenStudy (amistre64):

the "final" result is the construction of the plane equation itself; we need the normal vector as part of it

OpenStudy (anonymous):

Find an equation of the tangent plane to the surface z+ y/z = 1 at the point (4, 3, -1).